# This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible

### 20 Mar 2024 v1

@@ Abstract
Index Terms:
Automated Control Systems, Adversarial Attacks, Defense methods, Fault Detection and Diagnosis, Tennessee Eastman Process.
[…]
=-15pt
I Introduction
@@
The rest of the paper is organized as follows. Section II gives a review on attacks and defense approaches in the area. Section III describes methods of attacks and defenses used in our work. In section IV we describe experiments and discuss them in section V. Finally, section VI concludes the results.
II Review
@@
II-A Classification of attacks on machine learning models
@@
In evasion attacks (often called adversarial examples) [5, 6, 7], an attacker interacts with a trained machine learning model and manipulates its behavior by perturbing input samples during testing. The term ”evasion” implies that the attacker not only aims to cause the model to behave incorrectly but also seeks to evade detection by both human and automated defense mechanisms.
@@
II-B Most common methods of attack and defense
@@ II-C Summary of the review
III Methods
III-A Fault diagnosis methods
@@ III-B1 Random noise
III-B2 Fast Gradient Sign Method (FGSM)
@@ III-B3 FGSM distillation
III-B4 Projected Gradient Descent (PGD)
@@ IV Experiments
The description of our experiments is divided into 4 subsections. The FDD models subsection describes the training process of neural networks with different architectures. The next subsection shows how the accuracy of the models changes under different types of attack. Further, various methods for protecting models and their properties are shown. Finally, on the basis of the results of experiments, we also proposed and evaluated an approach consisting of a combination of two defense methods. All final results can be found in Fig. 7 and Tables VI-VIII.
IV-A FDD models
For our experiments, we used three models of neural networks with different architectures. To make the models differ from each other more, they contain different numbers of parameters and were trained for different numbers of epochs. The first model is a multilayer perceptron (MLP) consisting of two linear layers and containing 3 452 949 parameters. The second one is based on recurrent gated units (GRU) and containing 204 565 parameters. We also used temporal convolutional network (TCN) with 151 935 parameters. Data were standardized with a standard deviation of 1. Sliding window size was 32, which is a compromise between the accuracy of the models and the duration of the experiments. All models were trained on the TEP dataset for 20, 5 and 10 epochs for MLP, GRU and TCN respectively. The accuracy metrics for fault diagnosis on non-attacked data are presented in Table I. Combinations of the number of parameters and training epochs are selected on the validation set.
TABLE I: Accuracy of unprotected models on normal data.
@@ IV-B Attacks on FDD models
IV-C Protection of FDD models
All three neural network architectures have proven to be highly vulnerable to adversarial attacks and require protection methods. The defense methods studied in our research have many variations and parameters for selecting. It is not possible to conduct experiments for all combinations of settings and models in an adequate period of time. Therefore, we took only the TCN model, which has the fastest inference, to select more optimal settings for defense methods adjustment. Experiments conducted for each type of protection are described in following subsections. After setting up, the defense methods were applied to all FDD models and the final results are presented in Fig. 7 and Tables VI-VIII.
IV-C1 Adversarial training
@@
Adding more different perturbed data to the training process increases the average robustness of the model to adversarial attacks. But the quality with normal data decreases. Table II shows the accuracy of the TCN model on non-attacked data before and after adversarial training. Training with PGD adversarial samples showed better robustness from all attack types but worse quality in non-attacked mode. We used this setting for the final comparison of all defense methods.
TABLE II: Accuracy of the TCN model protected by adversarial training on normal data.
@@ IV-C2 Autoencoders
Accuracy metrics on non-attacked data can be seen in Table III. It is significantly lower than on unprotected model, but there is an opportunity to experiment with advanced autoencoder architectures in further research.
TABLE III: Accuracy of the TCN model protected by autoencoder on normal data.
@@ IV-C3 Quantization
Table IV shows the accuracy of the TCN model protected by quantization with different numbers of discrete values on non-attacked data.
TABLE IV: Accuracy of the TCN model protected by quantization with different parameter $n$ on normal data.
@@ IV-C5 Distillation
IV-D Combination of defense methods
@@
Quantization defense method having $2^{8}$ discrete values is more vulnerable to adversarial attacks than the one having $2^{5}$ values. But its combination with adversarial training gives partially better results, especially on normal, non-attacked data (Table V).
TABLE V: Accuracy of the TCN model protected by combination of adversarial training and quantization defense method on normal data.
@@ Appendix A Additional tables and figures
In this section we present the final comparison of all combinations of models, attacks and defenses (Tables VI, VII, VIII and Fig. 7).
@@
TABLE VI: Accuracy of protected and unprotected MLP model after adversarial attacks with different $\epsilon$ values.
@@
TABLE VII: Accuracy of protected and unprotected GRU model after adversarial attacks with different $\epsilon$ values.
@@
TABLE VIII: Accuracy of protected and unprotected TCN model after adversarial attacks with different $\epsilon$ values.

### 21 Mar 2024 v2

# Adversarial Attacks and Defenses in Automated Control Systems: A Comprehensive Benchmark

@@ Abstract
Keywords Automated Control Systems, Adversarial Attacks, Defense methods, Fault Detection and Diagnosis, Tennessee Eastman Process.
1 Introduction
@@
The rest of the paper is organized as follows. Section 2 gives a review on attacks and defense approaches in the area. Section 3 describes methods of attacks and defenses used in our work. In section 4 we describe experiments and discuss them in section 5. Finally, section 6 concludes the results.
2 Review
@@
2.1 Classification of attacks on machine learning models
@@
In evasion attacks (often called adversarial examples) [5, 6, 7], an attacker interacts with a trained machine learning model and manipulates its behavior by perturbing input samples during testing. The term "evasion" implies that the attacker not only aims to cause the model to behave incorrectly but also seeks to evade detection by both human and automated defense mechanisms.
@@
2.2 Most common methods of attack and defense
@@ 2.3 Summary of the review
3 Methods
3.1 Fault diagnosis methods
@@ 3.2.1 Random noise
3.2.2 Fast Gradient Sign Method (FGSM)
@@ 3.2.3 FGSM distillation
3.2.4 Projected Gradient Descent (PGD)
@@ 4 Experiments
The description of our experiments is divided into 4 subsections. The FDD models subsection describes the training process of neural networks with different architectures. The next subsection shows how the accuracy of the models changes under different types of attack. Further, various methods for protecting models and their properties are shown. Finally, on the basis of the results of experiments, we also proposed and evaluated an approach consisting of a combination of two defense methods. All final results can be found in Fig. 7 and Tables 6-8.
4.1 FDD models
For our experiments, we used three models of neural networks with different architectures. To make the models differ from each other more, they contain different numbers of parameters and were trained for different numbers of epochs. The first model is a multilayer perceptron (MLP) consisting of two linear layers and containing 3 452 949 parameters. The second one is based on recurrent gated units (GRU) and containing 204 565 parameters. We also used temporal convolutional network (TCN) with 151 935 parameters. Data were standardized with a standard deviation of 1. Sliding window size was 32, which is a compromise between the accuracy of the models and the duration of the experiments. All models were trained on the TEP dataset for 20, 5 and 10 epochs for MLP, GRU and TCN respectively. The accuracy metrics for fault diagnosis on non-attacked data are presented in Table 1. Combinations of the number of parameters and training epochs are selected on the validation set.
Table 1: Accuracy of unprotected models on normal data.
@@ 4.2 Attacks on FDD models
4.3 Protection of FDD models
All three neural network architectures have proven to be highly vulnerable to adversarial attacks and require protection methods. The defense methods studied in our research have many variations and parameters for selecting. It is not possible to conduct experiments for all combinations of settings and models in an adequate period of time. Therefore, we took only the TCN model, which has the fastest inference, to select more optimal settings for defense methods adjustment. Experiments conducted for each type of protection are described in following subsections. After setting up, the defense methods were applied to all FDD models and the final results are presented in Fig. 7 and Tables 6-8.
4.3.1 Adversarial training
@@
Adding more different perturbed data to the training process increases the average robustness of the model to adversarial attacks. But the quality with normal data decreases. Table 2 shows the accuracy of the TCN model on non-attacked data before and after adversarial training. Training with PGD adversarial samples showed better robustness from all attack types but worse quality in non-attacked mode. We used this setting for the final comparison of all defense methods.
Table 2: Accuracy of the TCN model protected by adversarial training on normal data.
@@ 4.3.2 Autoencoders
Accuracy metrics on non-attacked data can be seen in Table 3. It is significantly lower than on unprotected model, but there is an opportunity to experiment with advanced autoencoder architectures in further research.
Table 3: Accuracy of the TCN model protected by autoencoder on normal data.
@@ 4.3.3 Quantization
Table 4 shows the accuracy of the TCN model protected by quantization with different numbers of discrete values on non-attacked data.
Table 4: Accuracy of the TCN model protected by quantization with different parameter $n$ on normal data.
@@ 4.3.5 Distillation
4.4 Combination of defense methods
@@
Quantization defense method having $2^{8}$ discrete values is more vulnerable to adversarial attacks than the one having $2^{5}$ values. But its combination with adversarial training gives partially better results, especially on normal, non-attacked data (Table 5).
Table 5: Accuracy of the TCN model protected by combination of adversarial training and quantization defense method on normal data.
@@ Appendix A Additional tables and figures
In this section we present the final comparison of all combinations of models, attacks and defenses (Tables 6, 7, 8 and Fig. 7).
@@
Table 6: Accuracy of protected and unprotected MLP model after adversarial attacks with different $\epsilon$ values.
@@
Table 7: Accuracy of protected and unprotected GRU model after adversarial attacks with different $\epsilon$ values.
@@
Table 8: Accuracy of protected and unprotected TCN model after adversarial attacks with different $\epsilon$ values.